40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 20 23:55:51 2019, gautam, Update, ALS, ALS delay line electronics delayLineDemod.pdf.zip
    Reply  Thu Mar 7 01:06:38 2019, gautam, Update, ALS, ALS delay line electronics 
       Reply  Tue Mar 12 22:51:25 2019, gautam, Update, ALS, ALS delay line electronics 
       Reply  Wed Mar 13 01:27:30 2019, gautam, Update, ALS, ALS delay line electronics DFDcal.pdfDFDnoise.pdf
          Reply  Sun Mar 17 00:42:20 2019, gautam, Update, ALS, NF1611 cannot be shot-noise limited? PDnoise.pdf
             Reply  Thu Mar 28 19:40:02 2019, gautam, Update, ALS, BeatMouth with NF1611s assembled IMG_7381.JPGIMG_7382.JPGrelTF_schem.pdfdarkNoise.pdf
                Reply  Fri Mar 29 21:00:06 2019, gautam, Update, ALS, BeatMouth with NF1611s installed IMG_7384.JPGIMG_7385.JPG
                   Reply  Sun Mar 31 15:05:53 2019, gautam, Update, ALS, Fiber beam-splitters not PM IMG_7384.JPG
                      Reply  Wed Apr 3 09:04:01 2019, gautam, Update, ALS, Note about new fiber couplers 
                         Reply  Wed Apr 3 12:32:33 2019, Koji, Update, ALS, Note about new fiber couplers 
                      Reply  Fri Apr 5 00:33:58 2019, gautam, Update, ALS, Promising IR ALS noise IMG_7388.JPGfreeSpace_IR_beat.pdf
                         Reply  Fri Apr 5 11:49:30 2019, gautam, Update, ALS, PSL + X green beat recovery 
                         Reply  Mon Apr 8 00:04:08 2019, gautam, Update, ALS, IR ALS noise budget DFDnoise.pdf
                            Reply  Mon Apr 8 18:28:25 2019, gautam, Update, ALS, EX Green PDH checkout IMG_7393.JPGgreenModeMatch.pdf
                               Reply  Mon Apr 8 23:52:09 2019, gautam, Update, ALS, EX Green PDH error monitor calibrated errMonCalib.pdferrMon.pngOLTF.pdfEX_frequencyNoises.pdf
                                  Reply  Tue Apr 9 00:16:22 2019, rana, Update, ALS, EX Green PDH error monitor calibrated 
                                     Reply  Tue Apr 9 00:18:19 2019, gautam, Update, ALS, EX Green PDH error monitor calibrated 
                                  Reply  Tue Apr 9 18:44:00 2019, gautam, Update, ALS, EX Green PDH discriminant measurement cavityFlashes.pdfmodelPDH.pdf
                            Reply  Tue Apr 9 19:07:12 2019, gautam, Update, ALS, IR ALS noise budget ALS_noiseBudget.pdfALS_noiseBudget.pdf
                               Reply  Thu Apr 11 01:10:05 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X EX_PDH_2kNoise.pdf
                                  Reply  Wed Apr 17 00:50:17 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X no more EX_PDHnoise.pdf
                                     Reply  Wed Apr 17 11:01:49 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X no more EX_PDH_specGram.pdf
                      Reply  Wed May 29 18:13:25 2019, gautam, Update, ALS, Fiber beam-splitters are now PM 
                         Reply  Fri May 31 15:55:16 2019, gautam, Update, ALS, PSL + X beat restored 
Message ID: 14478     Entry time: Wed Mar 13 01:27:30 2019     In reply to: 14475     Reply to this: 14480
Author: gautam 
Type: Update 
Category: ALS 
Subject: ALS delay line electronics 

This test was done, and I determine the frequency discriminant to be \approx 5 \mu \mathrm{V}/\mathrm{Hz} (for an RF signal level of ~2 dBm). 

Attachment #1: Measured and predicted value of the DFD discriminant for a few RF signal levels.

  • Methodology was to drive an FM (deviation = 25 Hz, fMod = 221 Hz, fCarrier ~ 40 MHz) with the Marconi, and look at the IF spectrum peak height on a SR785
  • The "Design" curve is calculated using the circuit parameters, assuming 4dB conversion loss in the mixer itself, and 3dB insertion loss due to various impedance matching transformers and couplers in the RF signal chain. I fudged the insertion/convertion loss numbers to get this curve to line up with the measurements (by eye).
  • For the measurement, I assume the value for FM deviation displayed on the Marconi is an RMS value (this is the best I can gather from the manual). I'll double checking by looking at the RFmon spectrum directly on the Agilent NA.
  • X axis calibrated by reading off from the RF power monitor using a DMM and using the calibration data from the bench.
  • I could never get the ratio of peak heights in Ichan/Qchan (or the other way around) to better than ~ 1/8 (by moving the carrier frequency around). Not sure I can explain that - small non-orthogonality between I and Q channels cannot explain this level of leakage.

Attachment #2: Measured noise spectrum in the 1Y2 (LSC) electronics rack, calibrated to Hz/rtHz using the discriminant from Attachment #1.

  • Something funky with the I channel for X, I'll re-take that spectrum.

I'm still waiting on some parts for the new BeatMouth before giving the whole system a whirl. In the meantime, I'll work on the EX and EY green setups, to try and improve the mode-matching and better characterize the expected suppressed frequency noise of the end NPROs - the goal here is to rule out the excess low-frequency noise that was seen in the ALS signals coming from unsuppressed frequency noise.

Bottom lines: 

  1. The DFD noise is at the level of ~ 10mHz/rtHz above 10 Hz. This justifies the need for whitening before ADC-ing.
  2. The measured signal/noise levels in the DFD chain are in good agreement with the "expected" levels from circuit component values and typical insertion/conversion loss values.
  3. Why are there so many 60 Hz harmonics???
Attachment 1: DFDcal.pdf  21 kB  Uploaded Wed Mar 13 02:29:46 2019  | Hide | Hide all
DFDcal.pdf
Attachment 2: DFDnoise.pdf  88 kB  Uploaded Wed Mar 13 02:29:53 2019  | Hide | Hide all
DFDnoise.pdf
ELOG V3.1.3-